# Advancing Image Super-resolution Techniques in Remote Sensing: A Comprehensive Survey

### v2

@@ 4.2.4 Zero-shot Learning
Zero-shot learning-based RSISR method only uses the intrinsic features of the image to build a training set for super-resolution reconstruction. Yang and Wu [318] proposed an enhanced zero-shot super-resolution method, which generates enhanced images and builds training sets through a content adaptive resampler (CAR) network. In addition, they introduced a convolutional block attention module (CBAM) and a residual module, which significantly improved the quality of image detail reconstruction. In [319], the authors proposed a zero-shot RSISR framework based on self tessellations and cascaded attention sharing mechanism, which achieves high-quality reconstruction by exploiting the structural continuity of remote sensing images without pairing HR images. Self-FuseNet proposed by Mishra and Hadar [90] adopts a forward generation paradigm and accelerates image reconstruction through a UNet architecture with wide skip connections, demonstrating strong generalization capabilities in RSISR. Cha et al. [91] proposed a meta-learning based zero-shot remote sensing image super-resolution (ZSSR) method, which regards multiple LR image sets as a set of ZSSR tasks to learn general super-resolution meta-knowledge, significantly reducing the computational cost of large-scale image processing.
5 Recent Development Characteristics
In 2015, CNN was first applied to SR tasks, and the milestone work SRCNN [32, 33] was proposed. Further, msiSRCNN [148] was proposed in 2016 for RSISR. Since then, various CNN-based RSISR methods [53, 60, 62] have emerged. In addition, GANs [198] and Transformer [229] were introduced into the RSISR field in 2018 and 2021, respectively. These architectural models have been widely studied and used in the field of RSISR. In 2022, diffusion models [249] are also introduced into RSISR, and many diffusion models-based RSISR methods have been proposed. Another new architecture is Mammba, which was introduced into RSISR in 2024. We anticipate that an increasing variety of deep learning models will be adopted for RSISR tasks, in order to further enhance performance.
5.2 Most RSISR methods are supervised
Most RSISR methods involving deep learning are supervised, because RSI usually has the characteristics of high geometric alignment accuracy, which means that paired training samples can be easily constructed by downsampling. In addition, mainstream evaluation metrics such as PSNR and SSIM require ground truth (i.e., real HR image) as a reference, which further demonstrates the advantages of supervised learning. Although unsupervised and self-supervised methods have emerged in recent years, they still face great limitations in the field of remote sensing, such as poor sample diversity and lack of transferable pre-training. Therefore, most RSISR methods still rely on the supervised paradigm to obtain stable reconstruction performance.
5.3 RSISR methods for different types of images
Most existing RSISR methods perform HR reconstruction on optical RSI, which is mainly collected from satellite images such as WorldView, Gaofen and Sentinel-2. Recently, some researchers started developing methods to conduct SR on different types of RSI. However, it is not uncommon to perform SR on different types of RSI. For example, in 2012, He et al. [20] proposed an SR method for SAR images. Specifically, they achieved SR reconstruction on a very small test set by combining multi-dictionary and sparse coding spatial pyramid machine training. Similar work includes RSMAP [63] and Guo et al. [113].
Recently, many SR methods for HSI and MSI have been proposed. These images contain rich spectral information in a large number of bands and have broad research prospects. For example, from 2021 to 2024, many HSISR methods were proposed, covering Transformer-based [78, 85], Diffusion-based [84, 93] and self-supervised learning methods [39, 74, 89]. This diversity of RSISR tasks reflects the characteristics of different types of images, such as noise in SAR, spectral fidelity in HSI, and temporal consistency in satellite images, which require customized solutions. However, it is desirable to develop a general SR framework for different types of RSI, which may face great challenges in model generalization, cross-modal knowledge transfer, and computational efficiency.
5.4 Research on benchmarks
Although RSISR has made significant progress in recent years, benchmark research for RSISR still faces challenges. Unlike many basic tasks in computer vision, there are few studies on RSISR benchmarks, and there is a lack of general and high-quality RSISR benchmark datasets, which makes model performance comparisons unstable. In addition, most existing benchmarks [320, 321] rely on idealized downsampling degradation models, which limits the real-world adaptability of the methods.
Until recently, some researchers have started to create benchmarks in the field of RSISR. For example, Kowaleczko1 er al. [322] proposed the first benchmark dataset MuS2 based on real multi-temporal LR images and HR reference images, which breaks through the limitations of traditional simulated data and enhances the practicality of RSISR. In addition, Wang et al. [323] constructed a real degradation benchmark dataset RRSISR based on spectral camera imaging to address the mismatch between simulated degradation and real scenes in RSISR, and proposed a reference table-based block exit (RPE) method to improve efficiency through dynamic calculation.
5.5 Application-oriented RSISR methods
Most existing RSISR methods do not consider downstream applications in the process of image reconstruction. Therefore, these RSISR methods generate visually-pleasing reconstructed images, but this may be suboptimal for downstream tasks. This leads to a growing gap between metric-based performance (e.g., PSNR, SSIM) and actual utility in downstream applications. Bridging this requires RSISR models aligned with task-specific objectives, not just perceptual quality.
In the past three years, some studies have start deeply coupling the SR process with specific remote sensing application tasks to achieve a paradigm shift from visual quality-oriented to application-oriented. For example, Zhang et al. [209] proposed an RSI reconstruction method based on SRGAN to construct a high-quality landslide training set from LR images. In addition, Kong et al. [220] proposed a dual RSS-GAN method to integrate high-resolution Planet Fusion data to spatially enhance NDVI and NIRv, significantly improving the temporal expression and estimation accuracy of the surface index. Recently, Sun et al. [234] proposed the RSISR framework based on Transformer and self-supervision mechanism to effectively improve the performance of image interpretation. Fan et al. [215] proposed a GAN-based RSISR method, which significantly improved the accuracy of vegetation index inversion and yield prediction by improving the spatial resolution of ginkgo tree RSIs. Compared with most RSISR methods, these methods directly consider the performance of downstream applications during the image reconstruction process, which provides reconstructed images that are more suitable for specific applications.
5.6 Programming frameworks
@@ 6.1 Available remote sensing image datasets
[…]
  • 1.
    UCMD (UC Merced Land-Use Dataset) [^3] http://weegee.vision.ucmerced.edu/datasets/landuse.html [324]. This dataset is an RSI dataset for land use research, which is obtained from USGS National Map Urban Area Imagery. Besides, it contains images of 21 different land classes, with 100 images in each class. The pixel resolution of the public domain images is 1 foot (0.3 meters), and the image pixel size is $256\times 256$.
  • 2.
@@
  • 5.
    KOSD (Kaggle Open Source Dataset) [^7] https://www.kaggle.com/c/draper-satellite-image-chronology [328]. The dataset contains more than 1,000 HR aerial photographs from Southern California, consisting of 350 training images and 1,370 test images.
  • 6.
@@
  • 8.
    DOTA [^10] https://captain-whu.github.io/BED4RS/ [331]. DOTA is a large-scale public dataset collected from taken by Google Earth, GF-2 and JL-1 satellites for aerial image object detection. It contains a total of 2,806 HR images with resolutions ranging from $800\times 800$ to $4000\times 4000$ pixels.
  • 9.
@@
  • 10.
    DFC2019 (IEEE Data Fusion Contest 2019) [^12] https://ieee-dataport.org/open-access/data-fusion-contest-2019-dfc2019 [333]. The dataset contains 2,783 multi-temporal HR satellite images taken by the WV-3 satellite as training sets and 50 test images. All images are provided in the form of regular tiles of $1024\times 1024$ pixels with a spatial resolution of sub-meter level.
  • 11.
    OPTIMAL-31131313https://drive.google.com/open?id=1Fk9a0DW8UyyQsR8dP2Qdakmr69NVBhq9 [334]. The dataset contains 1,860 images of 31 categories collected from Google Maps, each category consists of 60 images with a size of $256\times 256$ pixels.
  • 12.
    SIRI-WHU [^14] http://www.lmars.whu.edu.cn/prof_web/zhongyanfei/e-code.html [335]. The dataset includes 2,400 images covering 12 different categories of scenes, collected and produced by the RS-IDEA Research Group of Wuhan University (SIRI-WHU) from Google Earth, mainly covering urban areas in China. Each category contains 200 images, each with a size of $200\times 200$ pixels and a spatial resolution of 2m.
  • 13.
@@ 6.2 Evaluation metrics
In the field of RSISR, evaluation metrics are crucial for measuring the performance of algorithms such as reconstruction accuracy and fidelity, which provides an objective reference for the performance comparison of algorithms through quantitative calculation. In this section, we summarize the commonly used quantitative evaluation indicators for RSISR, and the details are as follows:
@@
where $l(I_{\textit{SR}},I_{\textit{HR}})$, $C(I_{\textit{SR}},I_{\textit{HR}})$ and $S(I_{\textit{SR}},I_{\textit{HR}})$ represent the similarity of luminance, contrast and structure respectively. In practice, SSIM is not sensitive enough to the high dynamic range (such as clouds, shadow areas) or complex textures (such as forests, farmlands) of remote sensing images. Therefore, some studies have proposed the mean structural similarity (MSSIM) and multi-scale structural similarity (MS-SSIM) to evaluate the super-resolution performance. A higher SSIM value implies that the super-resolution image has a higher similarity and better fidelity with the HR image.
@@
(4) Natural Image Quality Evaluator (NIQE): NIQE [339] is a no-reference image quality evaluation metric. This metric is based on natural image statistical modeling and measures the image visual quality by comparing the distance between the local statistical features of the image and the statistical distribution. The lower the NIQE value, the better the image visual quality.
@@
where $G(x,y)$ represents the gradient at the pixel $(x,y)$, and the denominator $(h-1)(w-1)$ is used for normalization. The larger the AG value, the more details and sharper edges the image contains, and generally has higher clarity.

### v3

@@ 4.2.4 Zero-shot Learning
Zero-shot learning-based RSISR method only uses the intrinsic features of the image to build a training set for super-resolution reconstruction. Yang and Wu [321] proposed an enhanced zero-shot super-resolution method, which generates enhanced images and builds training sets through a content adaptive resampler (CAR) network. In addition, they introduced a convolutional block attention module (CBAM) and a residual module, which significantly improved the quality of image detail reconstruction. In [322], the authors proposed a zero-shot RSISR framework based on self-tessellations and a cascaded attention sharing mechanism, which achieves high-quality reconstruction by exploiting the structural continuity of RSIs without pairing HR images. Self-FuseNet proposed by Mishra and Hadar [91] adopts a forward generation paradigm and accelerates image reconstruction through a UNet architecture with wide skip connections, demonstrating strong generalization capabilities in RSISR. Cha et al. [92] proposed a meta-learning-based zero-shot RSI super-resolution (ZSSR) method, which regards multiple LR image sets as a set of ZSSR tasks to learn general super-resolution meta-knowledge, significantly reducing the computational cost of large-scale image processing.
5 Modal-Specific RSISR Methods
Remote sensing modalities such as optical, SAR, infrared, hyperspectral, and LiDAR exhibit distinct sensing mechanisms and data characteristics. These inherent differences demand correspondingly tailored super-resolution strategies. For instance, optical imagery provides rich spectral and spatial textures but is susceptible to atmospheric conditions. Conversely, SAR data offers robustness against weather and lighting variations yet is characterized by inherent speckle noise and nonlinear distortions. Consequently, each modality presents unique RSISR challenges, necessitating specialized architectures and loss functions designed to leverage their respective data priors.
In this section, we systematically review RSISR methods for different remote sensing modalities, including optical, SAR, infrared, and lidar data, analyzing fusion strategies, performance gains, and open challenges to guide future advancements.
5.1 Optical RSISR Methods
Typically, optical RSIs exhibit rich spatial details, strong inter-band correlations, and complex degradations caused by atmospheric scattering and optical system aberrations. In recent years, super-resolution for optical RSIs has emerged as a modality-specific field, with innovations driven by the unique imaging mechanisms and statistical characteristics inherent to optical data. In this section, we perform a detailed survey on recent progress in optical RSISR methods.
Some studies leverage the intrinsic inter-band correlations in optical or multispectral data as a powerful prior to enable joint reconstruction. For example, Zhang et al. [323] introduced a pixel-wise attention residual network that adaptively fuses features from multiple bands, significantly enhancing spectral fidelity across Sentinel-2 and Landsat datasets. Building upon this, Wang et al. [324] proposed a multi-scale CNN-Transformer hybrid (MSCT), which extracts local spectral features via adaptive residual dense blocks and models inter-band dependencies through Transformer attention, showing improved generalization on both RGB and multispectral inputs. To address the frequent loss of high-frequency texture and edge details in optical super-resolution, several works introduced frequency-aware generative architectures. Zhao et al. [253] developed EHC-DMSR, a diffusion-based model that enforces high-frequency consistency using Fourier-based constraints, improving the reconstruction of fine structures. In addition, Zhang et al. [113] proposed a detailed recovery based on GANs (DRGAN) for RSISR, which leverages dynamic convolution and self-attention to enhance high-frequency textures and mitigate artifacts and over-smoothing. Meanwhile, Liu et al. [165] a multiscale residual dense network (MRDN) for RSISR, which reuses multi-scale residuals across layers, preserving subtle details from the original low-resolution inputs.
Considering the specific degradation issues of optical systems, such as point spread functions (PSF), sensor limitations, and atmospheric effects, researchers have introduced physically informed architectures for RSISR. For instance, Liu et al. [325] built DLANet based on a dual-resolution design aligned with Gaussian blur priors, simulating optical sampling effects through low-resolution spatial branches. Chen et al. [326] proposed RDAF-GAN for RSISR, which incorporated a compound degradation model (blur, noise, compression) to generate realistic LR samples, improving generalization in real-world conditions. Sun et al. [261] introduced a distillation-based Transformer-CNN hybrid network, where a teacher model captures long-range degradation cues (e.g., haze) to guide physically plausible reconstruction by a lightweight student network.
Given the deployment constraints of optical satellites and edge devices, computational efficiency has become a critical issue. Wang et al. [327] designed the Context Transformation Layer (CTL) as a lightweight substitute for conventional 3×3 convolutions, achieving competitive performance on UC Merced with reduced parameter counts. Chen et al. [251] adopted a residual Mamba module based on a state-space model to capture long-range dependencies with linear complexity, addressing the scalability bottleneck for large-scale optical imagery. In addition, Li et al. [244] proposed an aggregation connection transformer (ACT-SR), which introduces row-column decoupled attention to suppress visual artifacts in Transformer-based reconstruction, improving perceptual consistency in urban and agricultural regions. To enhance interpretability and structural fidelity, Shi et al. [328] proposed DMUNet, an ADMM-based dual-branch unfolding network with dedicated edge reconstruction branches and a lightweight multiscale fusion module.
In summary, modality-driven RSISR research demonstrates compelling advantages by integrating physics-aware architectures with the semantic foundations of optical data. Crucially, advances in harnessing cross-band correlations, embedding degradation physics, and optimizing edge deployment efficiency underscore how deep domain knowledge of optical sensing mechanisms enables designing more effective, interpretable, and hardware-compatible super-resolution systems.
5.2 SAR RSISR Methods
Synthetic Aperture Radar (SAR) images possess unique physical properties, such as scattering matrix diversity, elevation sparsity, and speckle-resolution coupling, which limit conventional imaging resolution. These properties manifest differently across modalities: polarimetric systems require multi-channel optimization, tomographic setups are constrained by narrow elevation apertures, and single-channel systems face an inherent tradeoff between speckle suppression and spatial resolution. Recent RSISR techniques exploit these physical structures as reconstruction priors, overcoming traditional radar limits and mitigating acquisition-induced constraints. This section presents a detailed review of the RSISR method based on SAR data modalities.
Early work by Pastina et al. [329] pioneered Polarimetric SAR (PolSAR) super-resolution by applying parametric spectral estimation methods to jointly process multi-polarization channels (HH, VV, HV, VH). More recently, some studies employed deep learning techniques to perform detailed reconstruction. For example, Lin et al.[330] introduced a CNN with residual compensation for full-polarization SAR SR. In addition, Dai et al. [199] proposed a Polarimetric Contexture Convolutional Network (PCCN) for PolSAR SR, which encodes the polarimetric-spatial cube into a matrix representation and employs dual-branch feature extraction with hierarchical fusion to enhance reconstruction fidelity. In Tomographic SAR (TomoSAR), the limited elevation resolution resulting from sparse baseline sampling significantly constrains accurate 3D reconstruction. To address this challenge, recent studies have explored inversion methods that enhance elevation resolution through super-resolution techniques. For example, Zhu et al. [331] framed TomoSAR inversion as a spectral analysis problem and proposed the compressive sensing (CS)-based SL1MMER algorithm, exploiting the sparsity of elevation signals for SR. Further, Zhu and Bamler [332] rigorously quantified the theoretical limits of SL1MMER in terms of localization accuracy, minimum separable scatterer distance, and required acquisition numbers for robust SR. In addition, Wu et al. [333] proposed an RSISR approach for MIMO array TomoSAR under low SNR and limited antenna elements, which combined CS-based spatial filtering with deep neural network regression for improved scatterer localization in the elevation direction.
On the other hand, He et al. [20] proposed a multi-dictionary compressive sensing method for SAR SR using sparse coding spatial pyramids. Building on CS foundations, Wei et al. [334] subsequently introduced a hybrid scattering model incorporating line-segment scatterers (LSS) and rectangular-plate scatterers (RPS), solved via Multi-component ADMM. Concurrently, generative approaches gained traction. Specifically, Wang et al. [335] developed an SRGAN-based network for reconstruction and noise suppression. For target-level applications, Lee and Lee [336] created a robust method combining CS-based scattering center extraction with super-resolved system impulse response (IRF-S). Meanwhile, solutions addressing speckle noise emerged, including Importance Sampling UKF [337] and Deep GANs [338]. Finally, Dong et al. [339] established a sub-band decomposition framework with cross-domain learning and a novel amplitude-phase evaluation system. Recently, utilizing co-registered optical imagery to guide SAR SR reconstruction has emerged as a powerful strategy. Li et al. [340] proposed the Optical-Guided Super-Resolution Network (OGSRN), which incorporates an Enhanced Residual Attention Module (ERAM) and optical-space feedback mechanism to guide large-scale SAR SR. Zhao et al. [341] further advanced this area by proposing the Hierarchical Selective Fusion Mamba Network (HSFMamba), which uses state-space models and cross-modal feature selection for joint optical-guided SAR SR and denoising, supported by a large-scale aligned dataset.
Increasing attention focuses on resolving coupled degradations (e.g., resolution loss, speckle noise, defocusing) and end-to-end systems. For example, Chen et al. [342] proposed a degradation-aware diffusion super-resolution model (DADSR) that leverages contrastive learning for degradation representation and adaptively handles coupled degradations. Meanwhile, Zhang et al. [343] introduced an integrated deep learning network (DLIN) achieving end-to-end super-resolution imaging and target recognition through a 2D deep unfolding network and Transformer architecture.
SAR RSISR methods have evolved from modality-specific spectral estimation to deep learning frameworks that leverage intrinsic physical structures as reconstruction priors. While modality-tailored approaches have significantly advanced single-task SR performance, current research mainly focuses on three critical frontiers: multimodal guidance, coupled degradation handling and end-to-end task integration. Future efforts should prioritize the physics-informed architectures, large-scale aligned multi-sensor datasets for complex environments, and unified metrics for coupled SR-denoising evaluation. Real-time onboard processing remains a key challenge, demanding lightweight yet physics-compatible models.
5.3 Infrared RSISR Methods
Infrared (IR) imaging often suffers from low spatial resolution and poor texture detail due to the inherent limitations of detectors, optical diffraction, and challenging environmental conditions. These challenges hinder essential tasks such as military reconnaissance and night vision surveillance. To this end, infrared RSISR has emerged as a key solution to overcome hardware-imposed limitations.
Many studies have used multimodal fusion to solve the texture scarcity problem of infrared images, which mainly integrates the complementary visible information of different modal data. For example, Yang et al. [141] proposed a Soft-assignment based Multiple Regression (SMR) method that utilizes visible-IR joint training and cluster-adaptive dictionaries to enable real-time high-fidelity reconstruction. In addition, Jiang et al. [344] proposed a sensor-fusion network that utilizes cross-attention transformers and hierarchical distillation to recover thermal details while preserving radiometric accuracy. Chen et al. [345] developed a physics-guided fusion method for infrared super-resolution images, which integrates spectral inference from visible imagery, radiometric modeling, and correction using measured infrared data. Huang et al. [346] proposed DASRGAN, a dual-adaptation framework for IR super-resolution, which integrates Texture-Oriented Adaptation (TOA) to enhance texture and Noise-Oriented Adaptation (NOA) to minimize noise transfer, achieving state-of-the-art performance across multiple benchmarks.
Some studies focus on sparsity-driven methods, using IR’s inherent sparsity for efficient recovery and deeper exploration. Mao et al. [347] proposed a CS reconstruction method, which utilizes analytical measurement matrices and differential sparsification to achieve hardware-friendly stability. Zhang et al. [348] proposed a hybrid CS-CNN framework that utilizes sparse-domain initialization and deep denoising to enhance high-frequency components. In addition, advanced deep architectures leverage specialized representation learning for formidable infrared super-resolution, enabling higher-resolution imaging. For instance, He et al. [349] proposed a cascaded multi-receptive field network that leverages stage-wise upscaling with macro-to-micro refinement, achieving 8× super-resolution using a compact architecture. Cao et al. [236] developed the CFMB-T model, leveraging wavelet frequency separation and MTF-embedded degradation priors to focus reconstruction on critical high-frequency bands. Zhu et al. [350] introduced a pseudo-texture transfer algorithm that employs cross-modal feature matching and replacement for semantically aligned visible texture injection. Dan et al. [351] designed the PIRN architecture with phase-invariant augmentation and deformable convolution kernels to enhance spatial feature robustness. Zhang et al. [352] proposed a closed-loop learning system using coupled downsampling-SR generators to exploit real-image distributions across supervision paradigms. Recently, an IRSRMamba framework proposed by Huang et al. [353] addressed Mamba’s spatial fragmentation through wavelet-modulated state-space blocks and semantic consistency constraints. In addition, Lu and Su [354] proposed a multi-source domain adaptation method that utilizes visible-guided enhancement with thermal flux conservation to quadruple Mars IR resolution.
5.4 LiDAR RSISR Methods
LiDAR (Light Detection and Ranging) is a pivotal 3D sensing technology employed in autonomous driving, robotics, and environmental monitoring. Due to constraints in sensor cost, size, and energy efficiency, the acquired point clouds are typically sparse and low-resolution, undermining the accuracy of downstream tasks. To address this, LiDAR Point Cloud Super-Resolution (RSISR) has become a critical research focus, aiming to enhance point density and spatial fidelity through algorithmic post-processing.
Shan et al. [355] leveraged simulation-trained CNNs to enhance LiDAR resolution by projecting sparse point clouds into 2D range images, employing Monte Carlo dropout to filter uncertain predictions. Tian et al. [356] introduced a training-free geometric approach that segments and reorganizes ground/non-ground points to generate high-resolution outputs, eliminating training dependencies. Chen et al. [357] developed SGSR-Net, which integrates structural-semantic attention mechanisms with guided Monte Carlo filtering to boost geometric fidelity and SLAM compatibility. In addition, Ramirez et al. [358] proposed a GAN-based ”Hyperheight Data Cube” framework for satellite LiDAR, demonstrating significant spatial resolution enhancement in real-world validation. Collectively, these works demonstrate the field’s evolution from image-based enhancement toward geometric-aware modeling, semantic guidance, and generative frameworks in LiDAR RSISR research.
@@ 6.1 Available remote sensing image datasets
RSI-CB 2020 Crowdsource 60,000
$128\times 128$
$256\times 256$
  • UCMD (UC Merced Land-Use Dataset) [^3] https://reurl.cc/Y3XOQX [359]. This dataset is an RSI dataset for land use research, which is obtained from USGS National Map Urban Area Imagery. Besides, it contains images of 21 different land classes, with 100 images in each class. The pixel resolution of the public domain images is 1 foot (0.3 meters), and the image pixel size is $256\times 256$.
@@
  • KOSD (Kaggle Open Source Dataset) [^7] https://reurl.cc/9nprxa [363]. The dataset contains more than 1,000 HR aerial photographs from Southern California, consisting of 350 training images and 1,370 test images.
@@
  • DOTA [^10] https://captain-whu.github.io/BED4RS/ [366]. DOTA is a large-scale public dataset collected by Google Earth, GF-2, and JL-1 satellites for aerial image object detection. It contains a total of 2,806 HR images with resolutions ranging from $800\times 800$ to $4000\times 4000$ pixels.
@@
  • DFC2019 (IEEE Data Fusion Contest 2019) [^12] https://reurl.cc/6q0a9y [368]. The dataset contains 2,783 multi-temporal HR satellite images taken by the WV-3 satellite as training sets and 50 test images. All images are provided in the form of regular tiles of $1024\times 1024$ pixels with a spatial resolution of sub-meter level.
  • OPTIMAL-31131313https://reurl.cc/lYeRn6 [369]. The dataset contains 1,860 images of 31 categories collected from Google Maps, each category consists of 60 images with a size of $256\times 256$ pixels.
  • SIRI-WHU [^14] https://reurl.cc/ekOEnM [370]. The dataset includes 2,400 images covering 12 different categories of scenes, collected and produced by the RS-IDEA Research Group of Wuhan University (SIRI-WHU) from Google Earth, mainly covering urban areas in China. Each category contains 200 images, each with a size of $200\times 200$ pixels and a spatial resolution of 2m.
@@ 6.2 Evaluation metrics
In the field of RSISR, evaluation metrics are crucial for measuring the performance of algorithms such as reconstruction accuracy and fidelity, which provide an objective reference for the performance comparison of algorithms through quantitative calculation. In this section, we summarize the commonly used quantitative evaluation indicators for RSISR, and the details are as follows:
@@
where $l(I_{\textit{SR}},I_{\textit{HR}})$, $C(I_{\textit{SR}},I_{\textit{HR}})$ and $S(I_{\textit{SR}},I_{\textit{HR}})$ represent the similarity of luminance, contrast and structure respectively. In practice, SSIM is not sensitive enough to the high dynamic range (such as clouds, shadow areas) or complex textures (such as forests, farmlands) of RSIs. Therefore, some studies have proposed the mean structural similarity (MSSIM) and multi-scale structural similarity (MS-SSIM) to evaluate the super-resolution performance. A higher SSIM value implies that the super-resolution image has a higher similarity and better fidelity with the HR image.
@@
(4) Natural Image Quality Evaluator (NIQE): NIQE [374] is a no-reference image quality evaluation metric. This metric is based on natural image statistical modeling and measures the visual quality by comparing the distance between the local statistical features of the image and the statistical distribution. The lower the NIQE value, the better the image visual quality.
@@
where $G(x,y)$ represents the gradient at the pixel $(x,y)$, and the denominator (h1)(w1)(h-1)(w-1) is used for normalization. The larger the AG value, the more details and sharper edges the image contains, and generally has higher clarity.